OpenStudy (anonymous):

Write the standard form of the equation of the line described. through: (2, 4), perp. to y=-2/7x-5

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

y-4 = 7/2 (x-2) slope perp is negative reciprocal y = 7/2 x - 3

OpenStudy (anonymous):

well first you have to understand that for a line to be perpendicular to another the slope of such line must be the negative inverse of the given line. once your given that use point slope equation to find your line

OpenStudy (anonymous):

Didn't we do this already..?

OpenStudy (anonymous):

okay ill just stop haha

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

yes am sorry

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

you can continue

OpenStudy (anonymous):

The answer is above? U not understand it?

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

yes I understand

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

their result is 7x-2y=6

OpenStudy (anonymous):

It's the same....

OpenStudy (anonymous):

I used a different "standard form" but it's the same equation.

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

ok thank you very much

OpenStudy (anonymous):

y = 7/2 x - 3 Multiply by 2 is 2y = 7x -6 -> 2y -7x = 6.

OpenStudy (anonymous):

y = mx + b Standard form 1 ax + by = c Standard form 2

OpenStudy (anonymous):

I agree there should only be ONE standard form but there isn't....

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

hehehehhe

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

you do a god job

OpenStudy (anonymous):

OK :-) u r welcome.